OpenStudy (anonymous):

Juice-O-Rama is a refreshment stand specializing in a fruit drink mixture of orange juice and prune juice. The orange juice costs $1.20 per liter and the prune juice costs $1.40 per liter. If they need 8 liters of the final drink at a cost of $1.24 per liter, how many liters of each juice will they need to make their specialty drink?

OpenStudy (lgbasallote):

okay let x = number of liters of orange juice and y as number of liters of prune juice got it?

OpenStudy (anonymous):

okay

OpenStudy (lgbasallote):

it says the final product is 8 liters so x + y = 8

OpenStudy (lgbasallote):

then it says each liter of orange cost 1.2 dollars and each liter of prune costs 1.24 dollars then... 1.20 x + 1.24y = 1.40

OpenStudy (lgbasallote):

i believe the rest is systems of equations

OpenStudy (anonymous):

I understand now ilbasallote, thank you!!!!!!!

OpenStudy (anonymous):

Let oj and pj be equal to the final volume of orange juice and prune juice respectively. Solve the following for oj and pj:\[\text{Solve}[\{\text{oj}+\text{pj}=8,1.20 \text{ oj}+1.40\text{ pj}=8*1.24\},\{\text{oj},\text{pj}\}]\]{oj = 6.4 liters, pj = 1.6 liters}

OpenStudy (anonymous):

Thank you Robtobey!!

OpenStudy (anonymous):

Your are welcome.